OpenStudy (anonymous):

3/4g = 5/8 solve for g

OpenStudy (anonymous):

Is the question this \[\frac{ 3 }{ 4g } = \frac{ 5 }{ 8 }\] or \[\frac{ 3 }{ 4 }g = \frac{ 5 }{ 8 }\] ?

OpenStudy (anonymous):

the second example